40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11838     Entry time: Wed Dec 2 15:52:28 2015
Author: yutaro 
Type: Update 
Category: LSC 
Subject: restoration of POXDC 

I disconnected the cable that was connected to CH6 of the whitening filter in 1Y2, then connected POXDC cable to there (CH6). This channel is where POXDC used to connect.

Then I turned on the whitening filter for POXDC and POYDC (C1:LSC-POXDC FM1, C1:LSC-POYDC FM1) and changed the gain of analog whitening filter for POXDC and POYDC from 0 dB to 45 dB and from 0 dB to 39 dB, respectively (C1:LSC-POXDC_WhiteGain, C1:LSC-POYDC_WhiteGain).

ELOG V3.1.3-